# MobiFuse: Learning Universal Human Mobility Patterns through Cross-domain Data Fusion

### v1

@@ 4.1 Generative Data Fusion
Cross-Entropy Loss: $L_{CE}=-\sum_{c}w_{c}\cdot y_{c}\log(\hat{y}_{c})$ measures dissimilarity between predicted probabilities $\hat{y}_{c}$ and true labels $y_{c}$ with class weights $w_{c}$ to handle imbalanced classes; Transition Loss: $L_{TR}=-\frac{1}{N}\sum_{i=1}^{N}[t_{i}\log(\hat{t}_{i})+(1-t_{i})\log(1-\hat{t}_{i})]$ uses binary cross-entropy to penalize incorrect activity transitions, where $t_{i}$ and $\hat{t}_{i}$ are true and predicted transitions at time step $i$, ensuring coherent activity sequences; and Dynamic Time Warping Loss: $L_{DTW}=DTW(\hat{y},y)$ evaluates sequence similarity using DTW distance (Müller, 2007).
@@ 4.2.1 Supervised Transfer Learning for Knowledge Adaptation
When transfer to a region with a complete activity dataset (e.g., regional survey data), we employ progressive unfreezing as the transfer learning strategy. This approach balances knowledge preservation in the source domain with adaptation to the target domain, tuning three model components: embedding layers, Transformer layers, and an MLP, as illustrated in Figure 3. Following (Yosinski et al., 2014), input-near layers capture general features while output-near layers learn task-specific ones. By unfreezing the middle layers last (Peters et al., 2019), we mitigate catastrophic forgetting and facilitate fine-grained domain adaptation.
@@ 4.2.2 Semi-Supervised Transfer Learning for Mobility Pattern Adaptation
A supervised learning strategy is impractical for transferring the model to regions with only incomplete GPS data. To address this data sparsity issue, we develop a semi-supervised learning strategy as a key module to fuse survey data knowledge learned by the based model with the realism of GPS data in the target region.
@@
To avoid overfitting to the new synthetic data and ensure robustness, 20% of the previous iteration’s dataset is retained in each new training cycle. This retention strategy allows the model to refine its understanding by integrating information from both new synthetic data and a portion of the previously learned data, promoting stability and preventing drastic shifts in learned patterns. Furthermore, we differentiate real and synthetic data during training by applying a mask-based loss weighting technique. This method prioritizes real data in the loss computation, weighting it more heavily in the cross-entropy loss term as: $L_{real}=\frac{1}{N_{r}}\sum_{i=1}^{N}m_{i}\cdot L_{CE_{i}}$ and $L_{synthetic}=\frac{1}{N_{s}}\sum_{i=1}^{N}(1-m_{i})\cdot L_{CE_{i}}$, where $m_{i}$ is the mask (1 for real, 0 for synthetic), $N_{r}$ and $N_{s}$ are the numbers of real and synthetic points. Then the total loss function becomes: $L=w_{1}\cdot(w_{l}\cdot L_{real}+w_{s}\cdot L_{synthetic})+w_{2}\cdot L_{TR}+w_{3}\cdot L_{DTW}$.
4.2.3 Optimization Based Transfer Learning for Demographic Adaptation
@@
We propose a demographic adaptation mechanism through a learnable adapter matrix that modulates the activity reconstruction module’s output distribution. The adapter optimization is formulated as a optimization problem with reward-guided updates. We formulate a dual-criteria objective function that captures both global and local distribution alignment:$\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})=e^{-\alpha\cdot\text{RMSE}(\mathbf{P}_{g},\mathbf{D}_{g})}+\beta\cdot\text{KL}(\mathbf{P}_{g}\|\mathbf{D}_{g})$, where the exponential RMSE term ensures coarse distribution matching while the KL divergence term enforces fine-grained consistency between generated and target distributions. The adaptation weights are optimized through reward-weighted gradient descent, where $\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})$ serves as an adaptive scaling factor: $\mathbf{a}_{g}^{t+1}=\mathbf{a}_{g}^{t}+\eta\cdot\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})\cdot\nabla_{\mathbf{a}_{g}}\mathcal{L}$, with $\nabla_{\mathbf{a}_{g}}\mathcal{L}=(\mathbf{D}_{g}-\mathbf{P}_{g})$ providing the descent direction toward distribution matching.
To maintain numerical stability and interpretability, we impose bounded constraints on the adapter weights: $\mathbf{a}_{g}^{t+1}=\text{clip}(\mathbf{a}_{g}^{t+1},\epsilon,M)$, where $\eta$ denotes the learning rate, and $\epsilon,M$ define the feasible adaptation range. The optimization terminates upon reaching distribution convergence: $\|\mathbf{P}_{g}^{t}-\mathbf{D}_{g}\|_{2}\leq\delta\quad\forall g\in\mathcal{G}$, where $\delta$ controls the distribution matching precision. This formulation ensures stable convergence while maintaining interpretable demographic adaptation weights, which is elaborated at Algorithm 2.
@@ 4.3 Activity Location Assignment
To translate our generated activity chains into real-world use, we developed a location assignment module that merges spatial limits and traffic data, as in Figure 2 (4.3), which assigns each activity to specific TAZs through a layered approach: first fixing mandatory activities (home, work, school) based on commute patterns, then placing other activities while considering both distance spread and flow between fixed points. The process refines assignments until matching real spatial patterns that reflect regional traits. Through this assignment, our activity chains become real travel paths that can be tested against traffic data (detailed in Appendix A).
5 Experiments
We evaluate our framework through two case studies: Los Angeles (LA) County and Egypt. The LA study shows complete pipeline capabilities, while Egypt showcases transfer learning in data-scarce regions (as in Appendix C.2). Each component in our pipeline - semantic annotation, activity reconstruction, transfer learning (includes both activity pattern and demographic adaptation) - is essential in the data fusion process, as removing any component would prevent the generation of complete, demographically representative mobility patterns. We compare the output synthetic data against multiple sources, such as HTS, TUS and raw GPS points, as our baseline, which helped to reflect the disagreement on same measurement across different data sources.
5.1 Experimental Setup
Our dataset selection leverages the complementary strengths of different data sources: GPS trajectories provide large-scale mobility patterns but lack semantic context, which we complement with OpenStreetMap POI data. For activity pattern learning, we use HTS travel diary data as it contains complete activity chains, though its limited availability leads us to use the more accessible TUS data for demographic pattern modeling. This choice enhances our framework’s applicability across different regions. For validation, we select Caltrans PeMS traffic flow data and OpenStreetMap road network, as they provide comprehensive ground truth for assessing the realism of our generated mobility patterns.
We employ Jensen-Shannon Divergence (JSD) as our primary metric to evaluate distributional similarities across multiple dimensions: activity chain length, duration, type, and temporal patterns (start/end times). Lower JSD values indicate better alignment. For spatial validation, we use cosine similarity for Origin-Destination matrices and Mean Absolute Percentage Error (MAPE) for traffic metrics.
5.2 Model Performance Analysis
5.2.1 Base Model Evaluation on Survey Data
The performance of the base model $M^{N}$ in reconstructing 70% masked NHTS (National HTS) data and its transferability to the SCAG datasets is quantified in Table 2.
JSD Values Length Duration Type Start Time End Time
NHTS vs SCAG 0.017 0.009 0.059 0.018 0.013
$M^{N}$ on NHTS 0.017 0.002 0.001 0.001 0.005
$M^{N}$ on SCAG 0.021 0.003 0.017 0.004 0.007
$M^{S}$ on SCAG 0.011 0.004 0.004 0.001 0.001
The model achieves strong performance on NHTS data, with JSD values ranging from 0.001 to 0.017 across all metrics. However, directly applying the model to SCAG data results in performance degradation indicated by increasing JSD value, revealing significant regional differences in mobility patterns. The JSD values between NHTS and SCAG, which range from 0.009 to 0.059 across various activity attributes, quantify these differences and are visualized in Figure 4. Supervised transfer learning is applied to $M^{N}$ to obtain $M^{S}$, yielding low JSD values when reconstructing the SCAG dataset. This result shows the algorithm effectively adapts to different geographical contexts and data formats, despite initial disparities between SCAG and NHTS datasets.
5.2.2 Transfer Learning Evaluation on GPS and Survey Data
JSD values across iterations for the evaluation metrics are shown in Table 3, showing how the iterative transfer learning process improves alignment with the target region.
Table 3: Model evolution during iterative transfer learning
JSD Values Activity Type Start Time End Time
NHTS vs LA GPS 0.0508 0.0289 0.0367
Iteration_1 0.0116 0.0014 0.0037
Iteration_2 0.0033 0.0013 0.0038
Iteration_3 0.0022 0.0008 0.0027
Iteration_4 0.0025 0.0007 0.0019
Iteration_5 0.0046 0.0024 0.0052
Iteration_6 0.0077 0.0048 0.0081
In this table, only three metrics are included because the lack of ground truth in incomplete data renders activity duration and chain length unreliable for comparison, whereas the selected metrics are more accurate at the system level and can better reflect regional mobility patterns. Activity type JSD drops from 0.0508 at baseline to 0.0022 by Iteration 3. Similarly, start time and end time JSDs reach their lowest values, 0.0007 and 0.0019 respectively, at Iteration 4. Model performance deteriorates beyond Iterations 3 and 4, which appear to strike the best balance for reducing dissimilarities and adapting the model. This decline aligns with findings that AI models collapse when trained on recursively generated data (Shumailov et al., 2024), as over-reliance on synthetic data erodes nuanced or rare activities in the original data.
Group HTS vs Output TUS vs Output TUS vs HTS
Young M .083/.086 .008/.009 .014/.014
Young F .094/.085 .015/.013 .016/.011
Mid-aged M .048/.067 .018/.006 .016/.054
Mid-aged F .055/.066 .018/.007 .022/.059
Elder M .070/.066 .015/.009 .020/.055
Elder F .077/.071 .013/.009 .029/.063
Note: Values shown as (Worker/Non-worker); M: Male F: Female
Beyond activity durations, we evaluate temporal patterns across multiple scales: demographic groups (Figure 5), aggregate patterns (Figure 10), and activity-specific distributions (Figure 11). For demographic groups, the results show consistent patterns for mid-aged and elderly groups, while young groups show discrepancies due to limited young worker representation in HTS (2% of population). At the aggregate level, the model achieves strong alignment with both GPS data (JSD: 0.004) and NHTS patterns (JSD: 0.014), capturing key temporal features including peak periods. Major activities maintain their characteristic patterns through the transfer learning process, while our framework also captures early morning activities (before 6 AM) missing from HTS data due to survey limitations like underreporting and rounding effects (Bricka et al., 2012). This validates our approach’s ability to balance GPS and survey data while preserving demographic characteristics.
Figure 5: Temporal performance evaluation per demographic group based. Note: HTS lacks a Young Worker group; comparison for young individuals ignores employment status.
5.3 Traffic Simulation and Multi-Scale Validation of Synthetic Trajectories
To validate that our generated activity chains reflect realistic mobility patterns, we generated 1 million synthetic activity chains based on above methodology, and using location assignment module to assign perform a comprehensive real-world validation through large-scale traffic simulation, using MATSim (Horni et al., 2016) across LA County’s road network, enabling comparison with Caltrans traffic data (CalTrans, 2020). Our multi-scale validation shows strong performance across all spatial levels: at the LA regional level, the model achieves 0.997 cosine similarity for origin-destination matrices across eight sub-regions, demonstrating accurate reproduction of macro-level travel patterns and spatial interaction between regions; at the system level, MATSim simulations achieve MAPE of 4.97% for Vehicle Miles Traveled (VMT) and 1.16% for traffic speed, indicating that our synthetic populations generate realistic system-level traffic behavior across the entire network; at the corridor level, analysis of the I-405 corridor shows MAPE of 5.85-9.32% for traffic volume and 4.36-4.45% for speed across different directions, capturing local patterns including peak-hour congestion, validating the model’s ability to capture fine-grained local traffic patterns including peak-hour congestion. These results validate our framework’s ability to generate human travel diary with realistic mobility patterns, as evidenced by strong agreement with real-world traffic data (detailed in Appendix C.1.3).
6 Conclusion
This paper presents a framework that enables realistic mobility pattern generation through data fusion. Our approach addresses fundamental data challenges: semantically enriching and completing trajectories through LLM annotation and reconstruction, adapting to new regions through transfer learning, and incorporating demographic patterns without paired labels. We validate the generation based on real-world test case throughout large-scale simulation and showcase the strong performance. By fusing diverse data sources, our framework demonstrates the capability of overcoming individual data limitations. Future work could extend this framework to capture more complex social dynamics, including coordinated activities between individuals, large gathering events, and social network effects on mobility patterns, further enhancing the realism of generated trajectories.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning (ML), potentially benefiting transportation efficiency, urban development, and sustainability efforts. Our framework demonstrates how ML can harmonize diverse data sources to generate valuable insights while maintaining privacy. All data sources used are either publicly available (OpenStreetMap, census data, travel surveys) or commercially available in pre-anonymized form (GPS trajectories), with no collection of personal identifiers or sensitive information. The framework’s ability to work with anonymized data and generate synthetic patterns further ensures privacy preservation while enabling broad applicability for urban planning and policy making.
Appendix A Detailed Methodology of Activity Location Assignment
@@
The activity location assignment process transforms abstract activity chains into spatially-grounded trajectories by assigning each activity to specific TAZs. This process must balance multiple objectives: maintaining realistic commuting patterns, preserving feasible trip chains, and matching observed spatial distributions of activities across the region.
A.0.1 Mandatory Activity Assignment
@@
1. Distance Sampling: For each individual, we sample commute distances from demographic-specific distributions ($D_{md}$) derived from regional data. These distributions vary by sub-region to capture local commuting patterns, such as longer commutes in suburban areas versus shorter ones in urban cores.
2. Land Use Compatibility: Candidate TAZs are filtered based on land use compatibility. For work activities, we consider employment density and workplace location patterns; for school activities, we account for educational facility distribution.
3. Spatial Matching: The final TAZ selection optimizes for both the sampled distance and land-use suitability while ensuring the spatial distribution of mandatory activities matches regional patterns. This is formulated as: $Z_{md}=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\{d(h,z)-\hat{d}\}$ where $h$ is the home TAZ, $z$ is a candidate TAZ, and $\hat{d}$ is the sampled target distance.
A.0.2 Non-mandatory Activity Assignment
@@
1. Distance and Angular Constraints: Two key parameters guide the assignment: - Distance to next activity ($D_{nmd}$): sampled from activity-specific distributions - Angular deviation ($\theta$): measures the difference between the direct path to the next activity and the path to the next anchor point
2. Trip Chaining Optimization: Location selection considers both previous and next activities through a weighted objective function: $Z_{nmd}=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\{\alpha\cdot|d(z_{prev},z)-\hat{d}|+\beta\cdot|\theta(z_{prev},z,z_{next})-\hat{\theta}|\}$ where $\alpha$ and $\beta$ balance the importance of distance and directional constraints.
3. Temporal Feasibility: Assignments must ensure travel times between activities are realistic given the transportation network. This adds a constraint: $t(z_{i},z_{i+1})\leq T_{max}$ where $t(\cdot)$ is the travel time function and $T_{max}$ is the available time window.
A.0.3 Distribution Refinement

### v2

@@ 4.1.2 Data Augmentation by Activity Reconstruction
Cross-Entropy Loss is commonly used to measure the dissimilarity between predicted probability $\hat{y}_{c}$ and true activity labels $y_{c}$ for given class $c$ over the 96 time slots, as $L_{CE}=-\sum_{c}w_{c}\cdot y_{c}\log(\hat{y}_{c})$. The use of class weights $w_{c}$ balances the loss across potentially imbalanced activity classes.
Transition Loss uses binary cross-entropy to predict activity transitions by comparing predicted changes with true changes. It captures temporal dynamics by penalizing incorrect transitions, and encouraging accurate predictions at the correct time points. This approach helps generate realistic and coherent activity sequences, avoiding erratic jumps between activities.
$L_{TR}=-\frac{1}{N}\sum_{i=1}^{N}[t_{i}\log(\hat{t}_{i})+(1-t_{i})\log(1-\hat{t}_{i})]$ (3)
where N is the number of time steps, $t_{i}$ and $\hat{t}_{i}$ are true and predicted transitions at time step $i$ (1 if activity changed, 0 otherwise).
Dynamic Time Warping Loss (DTW), $L_{DTW}=DTW(\hat{y},y)$, is adopted to evaluate similarity between predicted sequence $\hat{y}$ and true sequence $y$ using DTW distance (Müller, 2007).
@@ 4.2.1 Supervised Transfer Learning for Knowledge Adaptation
When transferring to a region with a complete activity dataset (e.g., regional survey data), we employ progressive unfreezing as the transfer learning strategy. This approach balances knowledge preservation in the source domain with adaptation to the target domain, tuning three model components: embedding layers, Transformer layers, and an MLP, as illustrated in Figure 3. Following (Yosinski et al., 2014), input-near layers capture general features while output-near layers learn task-specific ones. By unfreezing the middle layers last (Peters et al., 2019), we mitigate catastrophic forgetting and facilitate fine-grained domain adaptation.
@@
Algorithm 1 Semi-Supervised Transfer Learning
  Input: Raw activity chain $A_{raw}$, initial model $M^{0}$
  Output: Synthetic activity chain $A_{syn}$, transferred model $M^{n}$
  Generate $A_{0}=M^{0}(A_{raw})$ and initialize $n=1$
  repeat
     Train $M^{n}$ using $A_{n-1}$ with masking
     Generate $A_{n}=M^{n}(A_{raw})$
     Compute JSD similarity $J=\text{JSD}(A_{raw},A_{n})$
     if $J$ has converged then
        $A_{syn}=A_{n}$
        break
     end if
     $n=n+1$
  until convergence or maximum iterations reached
@@ 4.2.2 Semi-Supervised Transfer Learning for Mobility Pattern Adaptation
A supervised learning strategy is impractical for transferring the model to regions with only incomplete GPS data. To address this data sparsity issue, we develop a semi-supervised learning strategy as a key module to fuse survey data knowledge learned by the base model with the realism of GPS data in the target region.
@@
$L_{real}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}m_{i}\cdot L_{CE_{i}}$ (4)
$L_{synthetic}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}(1-m_{i})\cdot L_{CE_{i}}$ (5)
where $m_{i}$ is the mask (1 for real, 0 for synthetic), $N_{r}$ and $N_{s}$ are the numbers of real and synthetic points, respectively.
The total loss function then becomes:
$L=w_{1}\cdot(w_{l}\cdot L_{real}+w_{s}\cdot L_{synthetic})+w_{2}\cdot L_{TR}+w_{3}\cdot L_{DTW}$ (6)
To avoid overfitting to the new synthetic data and ensure robustness, 20% of the previous iteration’s dataset is retained in each new training cycle. This retention strategy allows the model to refine its understanding by integrating information from both new synthetic data and a portion of the previously learned data, promoting stability and preventing drastic shifts in learned patterns.
Furthermore, we differentiate real and synthetic data during training by applying a mask-based loss weighting technique. This method prioritizes real data in the loss computation, weighting it more heavily in the cross-entropy loss term as:
where $w_{1}$, $w_{2}$, and $w_{3}$ are weights balancing the different loss components, while $w_{l}$ and $w_{s}$ control the relative importance of real versus synthetic data within the cross-entropy loss.
4.2.3 Optimization-Based Transfer Learning for Demographic Adaptation
@@
We propose a demographic adaptation mechanism through a learnable adapter matrix that modulates the activity reconstruction module’s output distribution. The adapter optimization is formulated as an optimization problem with reward-guided updates. We formulate a dual-criteria objective function that captures both global and local distribution alignment:
$\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})=e^{-\alpha\cdot\text{RMSE}(\mathbf{P}_{g},\mathbf{D}_{g})}+\beta\cdot\text{KL}(\mathbf{P}_{g}\|\mathbf{D}_{g})$ (7)
where the exponential RMSE term ensures coarse distribution matching while the KL divergence term enforces fine-grained consistency between generated and target distributions.
The adaptation weights are optimized through reward-weighted gradient descent, where $\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})$ serves as an adaptive scaling factor:
$\mathbf{a}_{g}^{t+1}=\mathbf{a}_{g}^{t}+\eta\cdot\mathcal{L}(\mathbf{P}_{g},\mathbf{D}_{g})\cdot\nabla_{\mathbf{a}_{g}}\mathcal{L}$ (8)
with $\nabla_{\mathbf{a}_{g}}\mathcal{L}=(\mathbf{D}_{g}-\mathbf{P}_{g})$ providing the descent direction toward distribution matching.
To maintain numerical stability and interpretability, we impose bounded constraints on the adapter weights:
$\mathbf{a}_{g}^{t+1}=\text{clip}(\mathbf{a}_{g}^{t+1},\epsilon,M)$ (9)
where $\eta$ denotes the learning rate, and $\epsilon,M$ define the feasible adaptation range. The optimization terminates upon reaching distribution convergence:
$\|\mathbf{P}_{g}^{t}-\mathbf{D}_{g}\|_{2}\leq\delta\quad\forall g\in\mathcal{G}$ (10)
where $\delta$ controls the distribution matching precision. This formulation ensures stable convergence while maintaining interpretable demographic adaptation weights, as shown in Algorithm 2.
  Output: Optimized adapter collection $\mathcal{A}$
  Process $\mathcal{T}$ to get target distributions $\{\mathbf{D}_{g}\}$
  Initialize $\mathcal{A}=\{\}$
  for $g\in\mathcal{G}$ do
     Initialize adapter weights $\mathbf{a}_{g}$
     repeat
        $\mathbf{Y}=f(\text{input})$
        $\mathbf{P}_{g}^{t}=\text{softmax}(\mathbf{Y}+\mathbf{a}_{g}^{t})$
        $\nabla_{\mathbf{a}_{g}}\mathcal{L}=\mathbf{D}_{g}-\mathbf{P}_{g}^{t}$
        $\mathbf{a}_{g}^{t+1}=\text{clip}(\mathbf{a}_{g}^{t+1},\epsilon,M)$
        if $\|\mathbf{P}_{g}^{t}-\mathbf{D}_{g}\|_{2}\leq\delta$ then
           break
        end if
     until $iteration=max\_iterations$
     $\mathcal{A}=\mathcal{A}\cup\{\mathbf{a}_{g}\}$
  end for
  return $\mathcal{A}$
@@ 4.3 Activity Location Assignment
To translate our generated activity chains into real-world use, we developed a location assignment module that merges spatial limits and traffic data, as in Figure 2 (4.3), which assigns each activity to specific TAZs through a layered approach: first fixing mandatory activities (home, work, school) based on commute patterns, then placing other activities while considering both distance spread and flow between fixed points. The process refines assignments until matching real spatial patterns that reflect regional traits.
@@
The activity location assignment process transforms abstract activity chains into spatially grounded trajectories by assigning each activity to specific TAZs. This process must balance multiple objectives: maintaining realistic commuting patterns, preserving feasible trip chains, and matching observed spatial distributions of activities across the region, as detailed in Algorithm 3.
4.3.1 Mandatory Activity Assignment
@@
Distance Sampling: For each individual, we sample commute distances from demographic-specific distributions ($D_{md}$) derived from regional data. These distributions vary by sub-region to capture local commuting patterns, such as longer commutes in suburban areas versus shorter ones in urban cores.
Land Use Compatibility: Candidate TAZs are filtered based on land use compatibility. For work activities, we consider employment density and workplace location patterns; for school activities, we account for educational facility distribution.
Spatial Matching: The final TAZ selection optimizes for both the sampled distance and land-use suitability while ensuring the spatial distribution of mandatory activities matches regional patterns. This is formulated as: $Z_{md}=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\{d(h,z)-\hat{d}\}$ where $h$ is the home TAZ, $z$ is a candidate TAZ, and $\hat{d}$ is the sampled target distance.
4.3.2 Non-mandatory Activity Assignment
@@
Distance and Angular Constraints: Two key parameters guide the assignment: - Distance to next activity ($D_{nmd}$): sampled from activity-specific distributions - Angular deviation ($\theta$): measures the difference between the direct path to the next activity and the path to the next anchor point
Trip Chaining Optimization: Location selection considers both previous and next activities through a weighted objective function: $Z_{nmd}=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\{\alpha\cdot|d(z_{prev},z)-\hat{d}|+\beta\cdot|\theta(z_{prev},z,z_{next})-\hat{\theta}|\}$ where $\alpha$ and $\beta$ balance the importance of distance and directional constraints.
Temporal Feasibility: Assignments must ensure travel times between activities are realistic given the transportation network. This adds a constraint: $t(z_{i},z_{i+1})\leq T_{max}$ where $t(\cdot)$ is the travel time function and $T_{max}$ is the available time window.
4.3.3 Distribution Refinement